# Unsupervised Detection of Anomalous Driving Patterns Using High Resolution Telematics Time Series Data

### v1

@@ 3.3 Adapting and Applying CTHMMs for Anomaly Detection in Telematics
$A_{ij,d}=\frac{\sum_{t}\mathbbm{1}(|z_{ijt,d}|\geq 3)}{T_{ij}}$
Note that the denominator is $T_{ij}$ only as forecast pseudo-residuals are not calculated at the first observation. These anomaly indices for response dimension $d$ across $N_{i}$ trips form an empirical distribution $\{A_{ij,d},j=1,...,N_{i}\}$ with empirical cumulative distribution function $F_{i,d}(p)$.
For group-level analysis, we will focus on the right tail of this distribution, reflecting extreme anomaly levels in dimension $d$. In particular, the maximum and $\alpha$th-percentiles will be of interest:
@@
    • The CTHMM should have sufficient (latent) states to capture the major combinations of response dimensions, which ensures accurate modelling of a driver’s (in individual-specific models) and a group’s (in a pooled model) normal driving behaviour. However, it is equally important to avoid an excessive number of states, which can lead to model overfitting and mistakenly fitting well to abnormal behaviour as well. The optimal number of states can be selected based on Akaike information criterion (AIC) and Bayesian information criterion (BIC).
@@
    • The forecast pseudo-residuals we consider for anomaly detection are based on probability integral transform, which applies to continuous distributions. While continuity adjustments can be made for discrete distributions, boundary points still pose problems. Harte, (2021) pointed out that pseudo-residuals are poor indicators of the model goodness of fit when observations are near or at the boundary of the domain. Similarly, we find that under these conditions, pseudo-residuals are also ineffective at detecting outliers. Although we can model all values of speed with a semi-continuous distribution (e.g., a mixture of a continuous distribution on the positive real line and a point mass at 0), the boundary point 0 is problematic for anomaly detection. For this reason, we exclude 0-speed observations by dividing each trip into intervals of non-zero speed and non-missing accelerations, capturing moments when the driver is actually driving and the vehicle is moving. Consequently, we assume these sub-intervals (instead of trips) of each driver are independent time series generated by the same model and the CTHMM is fitted to sub-intervals. For analysis and anomaly detection, we still evaluate on a trip basis: we first compute the pseudo-residuals for each sub-interval, then aggregate the total number of outliers across all sub-intervals to compute the anomaly index (outlier proportion) for the trip. By cutting trips into shorter, non-zero speed intervals, we aim to standardize the time series and reduce trip heterogeneity due to varying trip duration, road and weather condition, etc.
@@ 4.2.3 State-Dependent Distributions $\boldsymbol{\Theta}$
The maximization of $Q^{(m)}_{\boldsymbol{\Theta}}$ depends on the specific probability distributions that are employed. Generally speaking, one will have to compute the weighted MLE of the distribution parameters, with weights for the $u$-th latent state $\mathbb{E}(\mathbbm{1}_{\{Z_{ijl}=u\}}|\mathbf{Y}^{*},\mathbf{T}^{*},\boldsymbol{\Phi}^{(m-1)})$. In our proposed CTHMM, we assume the different dimensions $Y_{ijl,d}$ of the (multi-dimensional) response $\mathbf{Y}_{ijl}$ to be independent conditional on the latent state. Then conditional on the latent state, one can compute the (weighted) MLE of each dimension separately. In our Julia implementation, we have considered the following distributions: Gamma, Laplace, Log-Normal, Normal, von Mises, zero-inflated Gamma and zero-inflated Log-Normal. The MLE of the parameters of the selected distribution are provided in Appendix A.
@@ 5.2 Individual-Specific Model
We first test the ability of individual-specific models in identifying anomalous trips for each driver, with a particular focus on detecting the two aggressive trips. For each of the drivers D1 to D5, we fit a 20-state CTHMM to his/her seven trips. The choice of 20 states is made to balance between model flexibility and complexity. We have also tested a limited selection of 5, 10, 15, 20, and 25 states, with 20 states being optimal according to both AIC and BIC. This choice ensures that the model captures all major combinations of speed and accelerations without overfitting. More discussions on the robustness of the results with respect to the number of states are presented later in this subsection.
The state-dependent distributions are Gamma for speed, and Normal for both longitudinal and lateral accelerations. Initial guesses for parameter estimation are important in EM algorithm as they can influence the convergence of the algorithm and the final model fit, especially in the case when the algorithm converges to a local optimum instead of a global one. Parameter initializations of $\pi$ and $\mathbf{Q}$ are random, while state-dependent distribution parameters $\boldsymbol{\Theta}$ are initiated with k-means clustering. To robustify our analysis, we re-fit the model 100 times with different random initializations and analyze the average result.
@@ 6.1 Data Description
For both individual-specific and pooled models, telematics data undergo the same preprocessing steps as follows. First, we compute the longitudinal and lateral accelerations and align them with speed, as detailed in Section 2.2. Next, we divide each trip into intervals of non-zero speed and valid (non-missing) accelerations. For model training, trips must be at least 3 minutes long, and each trip sub-interval must contain at least 10 observations. This ensures that the CTHMM is trained on longer, informative sequences. For evaluation, the requirements are relaxed: trips must be at least 30 seconds long and contain at least 10 observations. These trips are similarly divided into intervals of non-zero speed and valid accelerations, with sub-intervals needing only at least 2 observations so that forecast pseudo-residuals can be computed; then all sub-intervals of a trip will be considered together to compute the anomaly index.
@@ 6.2 Individual-Specific Model
After model fitting, we compute the anomaly indices of each trip under its corresponding model. In Table 5, we present summary statistics of these indices for the 94 target trips and all other trips. We first observe that the target trips exhibit a slightly lower average anomaly index in speed compared to the other trips; however, this difference appears minimal as indicated by the other quartiles. As we learnt from the controlled study, speed alone does not fully capture driving behaviour and the analysis should focus on the accelerations. The summary statistics support this, revealing that the target trips have significantly higher anomaly indices in both acceleration components, with averages nearly 10 times higher. To further illustrate these differences, the left column of Figure 2(f) presents the empirical density plots of the anomaly indices for the two groups, with target trips shown in red and all other trips in blue. These plots help visualize the different distributional characteristics between the two groups, particularly in the acceleration components. Additionally, as we would like to rank the trips in each 3-day window by their level of anomalousness and ultimately identify the target trip, we can further consider the ‘normalized anomaly indices’, which is the ratio of each trip’s indices to the maximum indices within that window:
$\tilde{A}_{ij,d}=\frac{A_{ij,d}}{\max_{j}(A_{ij,d})}$
We present the summary statistics of these normalized indices in Table 6. The target trips have higher indices across all three dimensions compared to the other trips, with the differences much more pronounced for both acceleration components. This suggests that the target trips contain more outliers among the trips completed by the same driver, which will facilitate our identification of the target trips. It is important to note that we did not specify the target trips during model training nor pseudo-residuals computation, yet we observe distinct differences in anomaly indices and their normalized versions between the target and all other trips. The right column of Figure 2(f) displays the empirical density plots of the normalized anomaly indices for the two groups, with target trips represented in red and all other trips in blue.
@@
With the three raw or normalized anomaly indices as covariates, we consider three models to predict the probability that a trip is actually a target trip. The candidates are logistic regressions with and without cubic splines (i.e., Generalized Linear Model (GLM) and Generalized Additive Model (GAM), respectively) and XGBoost (eXtreme Gradient Boosting). We implement these models with standard R packages stats, mgcv and xgboost (with default parameter settings and 20 iteration rounds) respectively, and evaluate the area under the receiver operating characteristic (ROC) curve (AUC). ROC-AUC is a performance metric frequently used to evaluate the effectiveness for a binary classification model, which quantifies the model’s ability to distinguish between the two classes. A perfect model has an AUC of 1, while a model that randomly guesses has an AUC of 0.5. To robustify the performance evaluation, we perform 5-fold cross-validations (CV) for each set of covariates and model. In a $k$-fold cross validation, the dataset is split into $k$ subsets. In each fold, the model is trained on $k-1$ subsets, and the trained model is tested on the remaining subset. The $k$-fold CV then reports the average of the $k$ results. Our dataset is randomly partitioned into 5 subsets such that: first, all trips from a window are in the same subset; and second, the proportion of target trips are close (0.05234, 0.05263, 0.05882, 0.05740 and 0.04651).
@@
Finally, we fit the models to the whole dataset, and report the model summary of GAM in 9, while model summary of GLM and feature importance of XGBoost are provided in Appendix C. The findings show that outliers in both longitudinal and lateral accelerations are more significant predictors of target trips than outliers in speed, which is in line with observations from the controlled study.
@@
In our rental-car dataset, recall that our collaborator’s collision detection method produces accident triggers for only around 10% of claims. By applying the proposed method, we can effectively pinpoint identify target trips for investigation in the remaining 90% of cases. Moreover, leveraging the analysis of pseudo-residual points, as detailed in Section 5.2, allows us to locate each outlier and delve deeper into its nature. This facilitates the identification of underlying causes, such as harsh braking or cornering events relative to each driver’s driving patterns, and the provision of individualized feedback to mitigate such actions and improve driving practices.

### v2

@@ 3.3 Adapting and Applying CTHMMs for Anomaly Detection in Telematics
$A_{ij,d}=\frac{\sum_{t}\mathbbm{1}(|z_{ijt,d}|\geq 3)}{T_{ij}}$ (2)
Note that the denominator is $T_{ij}$ only as forecast pseudo-residuals are not calculated at the first observation. To rank each driver’s trips by their level of anomalousness, we further define the normalized anomaly index, which is the ratio of each trip’s index to the maximum index for that driver:
$\tilde{A}_{ij,d}=\frac{A_{ij,d}}{\max_{j}(A_{ij,d})}$ (3)
The (raw) anomaly indices for response dimension $d$ across $N_{i}$ trips form an empirical distribution $\{A_{ij,d},j=1,...,N_{i}\}$ with empirical cumulative distribution function $F_{i,d}(p)$. For group-level analysis, we will focus on the right tail of this distribution, reflecting extreme anomaly levels in dimension $d$. In particular, the maximum and $\alpha$th-percentiles will be of interest:
@@
    • The CTHMM should have sufficient (latent) states to capture the major combinations of response dimensions, which ensures accurate modelling of a driver’s (in individual-specific models) and a group’s (in a pooled model) normal driving behaviour. However, it is equally important to avoid an excessive number of states, which can lead to model overfitting and mistakenly fitting well to abnormal behaviour as well. As a general guideline, we suggest starting with a number of latent states greater than the number of response dimensions (e.g., at least 5 states for 3 response dimensions) and increasing the number of states incrementally. The optimal number of states can then be selected based on Akaike information criterion (AIC) or Bayesian information criterion (BIC).
@@
    • The forecast pseudo-residuals we consider for anomaly detection are based on probability integral transform, which applies to continuous distributions. While continuity adjustments can be made for discrete distributions, boundary points still pose problems. Harte, (2021) pointed out that pseudo-residuals are poor indicators of the model goodness of fit when observations are near or at the boundary of the domain. Similarly, we find that under these conditions, pseudo-residuals are also ineffective at detecting outliers. Although we can model all values of speed with a semi-continuous distribution (e.g., a mixture of a continuous distribution on the positive real line and a point mass at 0), the boundary point 0 is problematic for anomaly detection. For this reason, we exclude 0-speed observations by dividing each trip into intervals of non-zero speed and non-missing accelerations, capturing moments when the driver is actually driving and the vehicle is moving. Consequently, we assume these sub-intervals (instead of trips) of each driver are independent time series generated by the same model and the CTHMM is fitted to sub-intervals. For analysis and anomaly detection, we still evaluate on a trip basis: we first compute the pseudo-residuals for each sub-interval, then aggregate the total number of outliers across all sub-intervals to compute the anomaly index (outlier proportion) for the trip; see Figure 2 and the description that follows for an illustration. By cutting trips into shorter, non-zero speed intervals, we aim to standardize the time series and reduce trip heterogeneity due to varying trip duration, road and weather condition, etc.
    • Periods of low speed (near zero) may correspond to stops at traffic lights or other circumstances beyond the driver’s control, and therefore may not be relevant for identifying anomalous driving behaviour. While excluding 0-speed intervals was necessary for effective outlier detection, this adjustment also mitigates the concern about such low-speed periods. By focusing on non-zero speed intervals, representing moments when the driver is actively driving and the vehicle is moving, this approach ensures that stopping events, such as waiting at traffic lights, do not influence the anomaly detection process.
@@ 4.2.3 State-Dependent Distributions $\boldsymbol{\Theta}$
The maximization of $Q^{(m)}_{\boldsymbol{\Theta}}$ depends on the specific probability distributions that are employed. Generally speaking, one will have to compute the weighted MLE of the distribution parameters, with weights for the $u$-th latent state $\mathbb{E}(\mathbbm{1}_{\{Z_{ijl}=u\}}|\mathbf{Y}^{*},\mathbf{T}^{*},\boldsymbol{\Phi}^{(m-1)})$. In our proposed CTHMM, we assume the different dimensions $Y_{ijl,d}$ of the (multi-dimensional) response $\mathbf{Y}_{ijl}$ to be independent conditional on the latent state. Then conditional on the latent state, one can compute the (weighted) MLE of each dimension separately. In our Julia implementation, we incorporated the following distributions: Gamma, Laplace, Log-Normal, Normal, von Mises, zero-inflated Gamma, and zero-inflated Log-Normal. The MLE of the parameters for each selected distribution are provided in Appendix A of the Supplementary Material. The implementation is publicly available at https://github.com/ianwengchan/HMMToolkit.
4.3 Joint Modelling of Telematics and Computational Cost
In this work, we analyze multivariate time series of speed, longitudinal acceleration and lateral acceleration. Joint modelling of telematics responses provides a more complete picture of driving behaviour while also enhancing anomaly detection. Although it increases computational complexity, it provides a more comprehensive and robust risk assessment compared to analyzing each variable separately.
One key advantage of joint modelling is that it captures the interactions between different telematics responses, leading to a more complete representation of driving behaviour. Speed and acceleration complement each other, offering richer context than when analyzed individually. For instance, the same acceleration intensity at different speed ranges may indicate very different driving conditions: a harsh acceleration at low speeds could be a normal response in stop-and-go traffic, whereas the same acceleration at high speeds might signal aggressive driving.
Beyond improving behavioural modelling, joint modelling significantly enhances anomaly detection. As will be demonstrated in Section 5.2 (Tables 5 and 6), outlier identification depends not only on a single variable but on the combination of speed and acceleration components. Since pseudo-residual calculations and anomaly detection are influenced by which latent state is the most probable at each time point, and it depends on the other response dimensions as well, excluding key response dimensions weakens the model’s ability to identify true behavioural anomalies.
The trade-off between predictive accuracy and computational complexity primarily comes from an increase in the number of latent states. For example, if speed and longitudinal acceleration are each categorized into three levels (low, mid, high), a univariate CTHMM requires only 3 states, while a multivariate model may require 9 states. However, if both univariate and multivariate models are constrained to the same number of states, the multivariate model does not introduce significant additional computational costs. Under the assumption of conditional independence of response dimensions given a state, likelihood calculations remain computationally efficient as they only involve the multiplication of density functions across dimensions.
@@ 5.2 Individual-Specific Model
We first test the ability of individual-specific models in identifying anomalous trips for each driver, with a particular focus on detecting the two aggressive trips as a representative example. For each of the drivers D1 to D5, we fit a 20-state CTHMM to his/her seven trips. We tested a limited selection of models with 5, 10, 15, 20, and 25 states, with 20 states being optimal according to both AIC and BIC. This choice ensures that the model captures all major combinations of speed and accelerations without overfitting. More discussions on the robustness of the results with respect to the number of states are presented later in this subsection.
The state-dependent distributions are Gamma for speed, and Normal for both longitudinal and lateral accelerations. These choices are made based on the domain and characteristics of each response dimension. Gamma distribution is right-skewed and strictly positive, making it well-suited for modelling speed, which is often asymmetric with a long tail at higher values. Additionally, it has two parameters (shape and scale), allowing it to adapt to different degrees of skewness and dispersion in speed distributions. Normal distribution is appropriate for acceleration because acceleration values are often symmetrically distributed around zero, with both positive (acceleration) and negative (deceleration) values. It is also widely used due to its tractable properties and well-established estimation methods. In addition to optimizing the number of latent states, one can also explore alternative choices for state-dependent distributions. For example, a 20-state CTHMM could be fitted using either Lognormal or Gamma distributions for speed, with model selection guided by AIC or BIC. This flexibility allows for further refinement of the model based on empirical fit.
Initial guesses for parameter estimation are important in EM algorithm as they can influence the convergence of the algorithm and the final model fit, especially in the case when the algorithm converges to a local optimum instead of a global one. Parameter initializations of $\pi$ and $\mathbf{Q}$ are random, while state-dependent distribution parameters $\boldsymbol{\Theta}$ are initiated with k-means clustering. Users are encouraged to try multiple different initializations to reduce the chances of the model converging to a local optimum. To robustify our analysis, we re-fit the model 100 times with different random initializations and analyze the average result.
@@
While the anomaly index with the proposed outlier threshold of 3 standard deviations (SDs) provides strong anomaly detection results, we further examine the impact of threshold selection on performance. Table 4 presents the anomaly indices computed using alternative outlier thresholds of 2 and 4 SDs. Using a lower (higher) threshold increases (decreases) the anomaly index. We observe that with a 2-SD threshold, the two aggressive trips for each driver still exhibit higher anomaly indices in longitudinal acceleration and can be identified. However, with a 4-SD threshold, many indices drop close to zero, as it becomes highly unlikely for a normal forecast pseudo-residual to exceed 4 in absolute magnitude. As a result, the two aggressive trips can no longer be distinguished from other trips. Furthermore, while the 2-SD threshold still identifies the aggressive trips, it results in anomaly indices that are approximately 10 times higher than those computed with a 3-SD threshold (Table 3). This raises concerns about whether the trips are truly that anomalous and whether the anomaly index retains its intended meaning. For instance, under the 2-SD threshold, normal trips on the secondary road have an average of 3.84% outliers, whereas under the 3-SD threshold, this drops to 0.82%, which more accurately reflects them as following normal driving behaviour. These observations support the use of 3 standard deviations as the outlier threshold, as it effectively identifies anomalous trips and driving behaviour while maintaining interpretability.
Table 4: UAH-DriveSet: Alternative thresholds (2 and 4 standard deviations (SDs)) for calculating anomaly index in each dimension of telematics observations, computed from individual-specific CTHMMs with 20 states for each driver and averaged over 100 random initializations. Note: S stands for Secondary Road and M stands for Motorway; Z and Y are longitudinal and lateral accelerations, respectively.
2 SDs 4 SDs
Driver Trip Y_KF Z_KF Speed Y_KF Z_KF Speed
D1 S-Normal1 0.0740 0.0614 0.0126 0.0004 0.0000 0.0000
D1 S-Normal2 0.0690 0.0586 0.0038 0.0001 0.0000 0.0000
D1 S-Aggressive 0.0667 0.0675 0.0010 0.0008 0.0004 0.0000
D1 S-Drowsy 0.0657 0.0348 0.0100 0.0016 0.0021 0.0000
D1 M-Normal 0.0399 0.0389 0.0047 0.0001 0.0000 0.0000
D1 M-Aggressive 0.0565 0.0744 0.0327 0.0007 0.0011 0.0000
D1 M-Drowsy 0.0341 0.0294 0.0035 0.0000 0.0000 0.0000
D2 S-Normal1 0.0384 0.0185 0.0085 0.0000 0.0000 0.0000
D2 S-Normal2 0.0264 0.0299 0.0001 0.0008 0.0000 0.0000
D2 S-Aggressive 0.0413 0.1120 0.0318 0.0000 0.0000 0.0000
D2 S-Drowsy 0.1192 0.0431 0.0075 0.0025 0.0009 0.0000
D2 M-Normal 0.0253 0.0280 0.0086 0.0009 0.0000 0.0010
D2 M-Aggressive 0.0597 0.0999 0.0382 0.0012 0.0004 0.0000
D2 M-Drowsy 0.0894 0.0378 0.0060 0.0009 0.0004 0.0000
D3 S-Normal1 0.0349 0.0313 0.0028 0.0020 0.0000 0.0000
D3 S-Normal2 0.0341 0.0499 0.0042 0.0009 0.0000 0.0000
D3 S-Aggressive 0.0500 0.1163 0.0080 0.0009 0.0001 0.0000
D3 S-Drowsy 0.0728 0.0354 0.0024 0.0001 0.0000 0.0000
D3 M-Normal 0.0417 0.0384 0.0112 0.0002 0.0001 0.0000
D3 M-Aggressive 0.0528 0.0996 0.0528 0.0008 0.0007 0.0000
D3 M-Drowsy 0.0709 0.0144 0.0053 0.0010 0.0000 0.0000
D4 S-Normal1 0.0133 0.0433 0.0152 0.0000 0.0002 0.0000
D4 S-Normal2 0.0295 0.0344 0.0212 0.0000 0.0000 0.0000
D4 S-Aggressive 0.0612 0.1135 0.0126 0.0000 0.0017 0.0000
D4 S-Drowsy 0.0832 0.0436 0.0091 0.0016 0.0000 0.0000
D4 M-Normal 0.0310 0.0247 0.0145 0.0000 0.0000 0.0000
D4 M-Aggressive 0.0726 0.0943 0.0279 0.0011 0.0013 0.0000
D4 M-Drowsy 0.0657 0.0222 0.0055 0.0000 0.0000 0.0000
D5 S-Normal1 0.0220 0.0321 0.0032 0.0000 0.0000 0.0000
D5 S-Normal2 0.0504 0.0251 0.0105 0.0000 0.0000 0.0023
D5 S-Aggressive 0.0480 0.0884 0.0143 0.0004 0.0000 0.0000
D5 S-Drowsy 0.0655 0.0389 0.0053 0.0000 0.0000 0.0000
D5 M-Normal 0.0415 0.0551 0.0030 0.0000 0.0000 0.0000
D5 M-Aggressive 0.0823 0.0948 0.0055 0.0006 0.0001 0.0000
D5 M-Drowsy 0.0425 0.0399 0.0127 0.0006 0.0001 0.0000
@@ 6.1 Data Description
Figure 2: Speed time series (in $km/h$) of a sample trip. Shaded regions indicate intervals of non-zero speed, where all intervals are used for anomaly detection, but only the blue ones are used for model fitting.
For both individual-specific and pooled models, telematics data undergo the same preprocessing steps as follows. First, we compute the longitudinal and lateral accelerations and align them with speed, as detailed in Section 2.2. Next, we divide each trip into intervals of non-zero speed and valid (non-missing) accelerations, as shown by the shaded regions in Figure 2. For model training, trips must be at least 3 minutes long, and each trip sub-interval must contain at least 10 observations; these correspond to the blue shaded intervals in Figure 2. This ensures that the CTHMM is trained on longer, informative sequences. For evaluation, the requirements are relaxed: trips must be at least 30 seconds long and contain at least 10 observations. These trips are similarly divided into intervals of non-zero speed and valid accelerations, with sub-intervals needing only at least 2 observations so that forecast pseudo-residuals can be computed. Hence, pseudo-residuals are computed for all shaded intervals (both the blue and pink ones). Finally, all sub-intervals of a trip will be considered together to compute the anomaly index.
Remark: While we assume that all trips within each rental are performed by a single driver, it is possible for vehicles to have multiple drivers. In such cases, the model captures the dominant driving behaviour exhibited collectively by all drivers of the vehicle rather than individual-specific patterns. This can be viewed as a pooled model, which still accounts for vehicle-specific effects (e.g., horsepower) while incorporating driving patterns across multiple users. As demonstrated in Section 5.3, a pooled model can still effectively identify anomalous trips, even when multiple drivers contribute to the data.
A potential limitation arises when the dataset is imbalanced. For example, if one driver completes most trips and another completes only a single trip, that one-off trip — even if it is objectively safer — may still be classified as anomalous if it significantly deviates from the dominant driving pattern. However, this is an inherent characteristic of anomaly detection, and controlling for such cases would require additional contextual information. If information on driver changes is available, the model could be adapted simply by clustering trips by driver and training a separate model for each driver, thereby reducing the influence of mixed driving patterns.
@@ 6.2 Individual-Specific Model
After model fitting, we compute the anomaly indices of each trip under its corresponding model. In Table 8, we present summary statistics of these indices for the 94 target trips and all other trips. We first observe that the target trips exhibit a slightly lower average anomaly index in speed compared to the other trips; however, this difference appears minimal as indicated by the other quartiles. As we learnt from the controlled study, speed alone does not fully capture driving behaviour and the analysis should focus on the accelerations. The summary statistics support this, revealing that the target trips have significantly higher anomaly indices in both acceleration components, with averages nearly 10 times higher. To further illustrate these differences, the left column of Figure 3 presents the empirical density plots of the anomaly indices for the two groups, with target trips shown in red and all other trips in blue. These plots help visualize the different distributional characteristics between the two groups, particularly in the acceleration components. Additionally, as we would like to rank the trips in each 3-day window by their level of anomalousness and ultimately identify the target trip, we can further consider the normalized anomaly indices given in Equation 3, which is the ratio of each trip’s indices to the maximum indices within that window.
We present the summary statistics of these normalized indices in Table 9. The target trips have higher indices across all three dimensions compared to the other trips, with the differences much more pronounced for both acceleration components. This suggests that the target trips contain more outliers among the trips completed by the same driver, which will facilitate our identification of the target trips. It is important to note that we did not specify the target trips during model training nor pseudo-residuals computation, yet we observe distinct differences in anomaly indices and their normalized versions between the target and all other trips. The right column of Figure 3 displays the empirical density plots of the normalized anomaly indices for the two groups, with target trips represented in red and all other trips in blue.
@@
With the three raw or normalized anomaly indices as covariates, we consider three models to predict the probability that a trip is actually a target trip. The candidates are logistic regressions with and without cubic splines (i.e., Generalized Linear Model (GLM) and Generalized Additive Model (GAM), respectively) and XGBoost (eXtreme Gradient Boosting). We implement these models with standard R packages stats, mgcv and xgboost, respectively. For GAM, we specify a cubic regression spline (bs = ‘cr’), while for XGBoost, we use the package’s default parameter settings, adjusting only the number of iteration rounds to 20. This choice ensures reproducibility and consistency, as training log loss stabilizes around 0.001 at this point, indicating sufficient convergence without overfitting. While further optimization of the XGBoost model is possible, our primary focus is on demonstrating the usefulness of the proposed anomaly index in identifying accident-related trips.
We then evaluate model performance based on ROC-AUC (the area under the receiver operating characteristic curve). ROC-AUC is a performance metric frequently used to evaluate the effectiveness for a binary classification model, which quantifies the model’s ability to distinguish between the two classes. A perfect model has an AUC of 1, while a model that randomly guesses has an AUC of 0.5. To robustify the performance evaluation, we perform 5-fold cross-validations (CV) for each set of covariates and model. In a $k$-fold cross validation, the dataset is split into $k$ subsets. In each fold, the model is trained on $k-1$ subsets, and the trained model is tested on the remaining subset. The $k$-fold CV then reports the average of the $k$ results. Our dataset is randomly partitioned into 5 subsets such that: first, all trips from a window are in the same subset; and second, the proportion of target trips are close (0.05234, 0.05263, 0.05882, 0.05740 and 0.04651).
@@
Finally, we fit the models to the whole dataset, and report the model summary of GAM in Table 12, while model summary of GLM and feature importance of XGBoost are provided in Appendix C. The findings show that outliers in both longitudinal and lateral accelerations are more significant predictors of target trips than outliers in speed, which is in line with observations from the controlled study.
@@
In our rental-car dataset, recall that our collaborator’s collision detection method produces accident triggers for only around 10% of claims. By applying the proposed method, we can effectively pinpoint target trips for investigation in the remaining 90% of cases. Moreover, leveraging the analysis of pseudo-residual points, as detailed in Section 5.2, allows us to locate each outlier and delve deeper into its nature. This facilitates the identification of underlying causes, such as harsh braking or cornering events relative to each driver’s driving patterns, and the provision of individualized feedback to mitigate such actions and improve driving practices.